Please help me to solve 9/(x-5)-1=8/(x+5)
Thanks

Answer by Nate(3500)

9(x - 5)/(x - 5) - 1(x - 5) = 8(x - 5)/(x + 5)
9 + (-x + 5) = (8x - 40)/(x + 5)
9(x + 5) + (-x + 5)(x + 5) = (8x - 40)(x + 5)/(x + 5)
9x + 45 - x^2 + 25 = 8x - 40
9x + 70 - x^2 = 8x - 40
0 = x^2 - x - 110
110 = x^2 - x
1/4 + 440/4 = (x - 1/2)^2
441/4 = (x - 1/2)^2
+-21/2 = x - 1/2
1/2 +- 21/2 = x
11 = x and -10 = x